OpenStudy (anonymous):

Hello! Here is my question: I have in space a line AB where A (0,0,0) and B (3,5,10). I have a point C (4,7,9). I need to find on line AB a point D where the line CD intersects AB if this lines are perpendicular to each other. Please show step by step. Thanks.

OpenStudy (anonymous):

Find vector AB.

OpenStudy (amistre64):

ab = <3,5,10> right?

OpenStudy (amistre64):

you need to determine if the lines even intersect

OpenStudy (anonymous):

Very good. Now call D, xyz and find CD

OpenStudy (amistre64):

<3,5,10> <x, y, z> --------- 3x +5y +10z = 0

OpenStudy (amistre64):

3(x-4t) +5(y-7t) +10(z-9t) = 0 right?

OpenStudy (toxicsugar22):

chaguanas can u help me with a math question

OpenStudy (amistre64):

x = 3t y = 5t z = 10t is the equation of one line in space

OpenStudy (toxicsugar22):

amistre can u help me

OpenStudy (amistre64):

x = x0+4t y = y0+7t z = z0+9t is the other.... i think

OpenStudy (amistre64):

i can look at it, not sure if i can helptho :)

OpenStudy (anonymous):

I do not remember how to find th Perpendicular

OpenStudy (anonymous):

To find perpendicular is to find the normal, cross two vectors.

OpenStudy (anonymous):

Vector U*V=0 right?

OpenStudy (anonymous):

Yes, Amistre, found AB and CD. Actually, you just test them to see if they are perpendicular.

OpenStudy (anonymous):

AB*CD=0

OpenStudy (anonymous):

AB*CD=0

OpenStudy (anonymous):

Yes, if you dot them and get 0, they are perpendicular

OpenStudy (anonymous):

I will try. How do I find the normal, cross two vectors?

OpenStudy (anonymous):

Yes that is how you found the normal. But in this case by the way the question was asked you were led to the answer to your question without having to do that.

OpenStudy (anonymous):

Thanks